OpenStudy (anonymous):

A charge Q distributed uniformly along the edge of a ring of radius R. What is the electric potential due to this charge at a distance x=R from the center the ring along its axis?

OpenStudy (roadjester):

\[\huge V=k_e \int{dq \over r}=k_e\int{dq\over \sqrt{a^2+x^2} }\] \[\huge V={k_e \over{\sqrt{a^2+x^2}} }\int dq={k_eQ\over {\sqrt {a^2+x^2}}}\]